OpenStudy (anonymous):

Which expression represents the product of 6.5*10*4 and 2.4*10*3

OpenStudy (anonymous):

how do you get 1.56 instead of 15.6

OpenStudy (unklerhaukus):

\[6.5\times10^4 \times 2.4\times10^3\]\[=6.5 \times 2.4\times10^4 \times10^3\]\[=(6.5\times2.4)\times10^{4+3}\]\[=\cdots\]

OpenStudy (anonymous):

yeah i kno when u multiply 6.5X2.4 ull get 15.6 but the right answer is 1.56 how??

OpenStudy (unklerhaukus):

\[=15.6\times10^{4+3}\] \[=1.56\times10^{4+3+1}\]

OpenStudy (anonymous):

and how come the exponents are 8 instead of seven im soo confused please help i have regents after an hour

OpenStudy (anonymous):

but why did u change 15.6 into 1.56

OpenStudy (unklerhaukus):

why . because scientific notation has only one number before the decimal point,

OpenStudy (anonymous):

OMG thank yuhhh so in scientific notation theres always one number before the decimal point :D ?

OpenStudy (anonymous):

ALWAYS??

OpenStudy (unklerhaukus):

yeah,

OpenStudy (anonymous):

thank yuh then why do you add 1 to the exponent

OpenStudy (anonymous):

why isnt 10^7 ?

OpenStudy (anonymous):

??

OpenStudy (unklerhaukus):

\[=15.6\times10^{7}\]\[=\frac{15.6}{10}\times10\times 10^7\]\[=1.56\times10^{7+1}\]

OpenStudy (anonymous):

so you allways add one to the exponent

OpenStudy (unklerhaukus):

well if you had \[100.\] and you wanted to represent in scientific notation \[100.=100.\times1\]\[=100\times10^0=10^2\times10^0\]\[=1\times10^{2+0}=1\times10^2\]

OpenStudy (anonymous):

im sooo confused imma fail regents :((

OpenStudy (unklerhaukus):

or \[44.8=4.48\times10^1\] or\[5500.0=55.000\times10^3=5.5000\times10^2\]

OpenStudy (unklerhaukus):

each decimal place represents a factor of ten dec-i-mal means tenth

OpenStudy (anonymous):

ohhhhhh so ok can you try to give me problem and ill try to solve cuz i think i need a little practice

OpenStudy (unklerhaukus):

who would you write 501 in scientific notation

OpenStudy (unklerhaukus):

how*

OpenStudy (unklerhaukus):

there is an invisible decimal point directly right of the 1

OpenStudy (anonymous):

5.01X10^2

OpenStudy (unklerhaukus):

yep too easy huh?

OpenStudy (anonymous):

OMG yess you just saved my life

OpenStudy (anonymous):

then how do you write the scientific notation thing into regular form

OpenStudy (unklerhaukus):

\[5.002\times10^4=5.002\times10\times10\times10\times10=50020.\]

OpenStudy (anonymous):

then how do u get negative exponents?? and sorry if im wasting your time

OpenStudy (unklerhaukus):

write \[9.80532\times10^1+2\times 10^{-5}\]in regular notation

OpenStudy (unklerhaukus):

\[8\times10^{-2}=0.08\] \[3.2\times10^{-3}=0.0032\]

OpenStudy (unklerhaukus):

negative exponent means the decimal moves to the left .

OpenStudy (anonymous):

ohhhhhhhhhh

OpenStudy (anonymous):

can i ask another question

OpenStudy (unklerhaukus):

\[100=1\times10^2\]\[10=1\times10^1\]\[1=1\times10^0\]\[0.1=1\times10^{-1}\]\[0.01=1\times10^{-2}\]

OpenStudy (anonymous):

A company is designing a cylinder to hold marbles for a new game it is inventing. The cylinder has a height of 18 inches and a diameter of 6 inches. Find the volume of the cylinder to the nearest tenth of a cubic inch.

OpenStudy (unklerhaukus):

do you know the area of a circle?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is the formula V=Pi(r)^2h

OpenStudy (unklerhaukus):